OpenStudy (anonymous):

Solve for the indicated variable. A = 1 + prt; for r r=

OpenStudy (anonymous):

Same thing before A = 1+prt 1+prt = A -1 -1 prt = A-1 divide both sides by pt r = (A-1)/pt

OpenStudy (anonymous):

Do the same steps above, and you should be able to solve most of these.